Subject: 2004 GRImII status

From: Mark Klaene

Submitted: Tue, 20 Jan 2004 16:24:51 -0700

Message number: 733 (previous: 732, next: 734 up: Index)

GRIM II's single 256 x 256 NICMOS 3 chip is read out simultaneously with
four amplifiers, dividing the array into four quadrants, or "quads", which
can be considered semi-independently.  For several years, the generally
reliable character of GRIM has been intermittently (and randomly)
punctuated with instances where one or more of the readout quadrants fails
to read out properly.  The symptoms are either quads that return
inappropriate signal levels, or none at all, while adjacent quads function
normally.  The unpredictable quality of this problem meant that the
instrument would function with all four quads operating reliably for weeks
or months, then in one night we could potentially lose an hour or more
trying to recover when a quad would drop.  After some black magic, said
quad would return and GRIM would work reliably for the remainder of a

In November 2003 we saw the frequency of the occurrence of this problem
spike, to the extent that quadrants were dropped in nearly every image
GRIM returned.  At that point, GRIM was taken out of service while we
attempted to find a fix to the problem.  Dale Sanford (UC), one of 
GRIM's original architects, was brought in during December to 
evaluate GRIM's condition and to assess the nature of the problem. 
During his stay, the exterior (i.e., outside the dewar) electronics 
were tested for obvious faults, and we characterized the performance 
of what at the time were the still reliably functioning quads. Dale 
concluded that the root of the problem lies within the dewar. 
Although another annoying problem may be corrected with improved 
ventilation. This ventilation modification will be completed within 
the next week or so.  Any further repair work would require opening 
the dewar. An undertaking that will be very time consuming and with 
risks as there are a number of other repair and maintenance items 
that would have to be addressed once the dewar is opened.

Concurrently, the University of Colorado-built Near Infrared 
Camera/Fabry-Perot Spectrometer (NIC-FPS) is scheduled for completion 
and commissioning in mid-'04, with routine science operations to 
begin around the first of the new year.  The point of this message is 
to address how ARC near-IR programs will be affected for the 
remainder of 2004.

As of January 2004, GRIM exists in a semi-functional state.  Our
characterization of the three quadrants that were "alive" in December
indicates that the presence of a bad quad doesn't appear to affect the
adjacent quads.  We verified quantities such as gain, read noise,
linearity and dark current and found those values reasonably close to the
ones tabulated on the APO website  One somewhat consistent 
aspect of the problem is *which* quads tend to drop; i.e., the 
leftmost two quads, as viewed in IRAF, seem consistently alive, while 
the upper right quad is most flaky and the lower right quad somewhere 
in between.

The main impact of this problem on science programs is, again, the
unpredictability of the quadrants.  On a given night, all four may be
functional all night, one or two may drop either periodically or
permanently through a night, or some permutation of these possibilities
may exist.  Non-scripted, pointed observations would not suffer
unreasonably from this effect, but considering that most GRIM observations
are scripted, the current state of GRIM has a definite and significant
impact on the majority of GRIM science.  Observers considering 
proposing GRIM observations later in the year should consider these 
issues carefully.

Most importantly, all potential 2004 GRIM users should be prepared for the
circumstance that GRIM may be taken out of service either temporarily or
permanently, at any time and for any length of time, before NIC-FPS
becomes available.  The engineers at APO hope to keep GRIM running long
enough to overlap with early NIC-FPS service and to characterize the 
new instrument sufficiently to allow for correlation of its data with 
GRIM data, determination of zeropoint shifts, etc.  But this calendar 
year will be very tentative with respect to near-IR data and the 3.5m 
telescope.  Users should plan accordingly.  Also, any planned outages 
for service and repair will be announced to the community as far in 
advance as possible to keep the chaos to a minimum.

Observers faced with using GRIM in its current state should be on the
lookout for dropping quads during their observations, but there is no
evident cause to mistrust data the other functional quads are giving.  If
a functional quad drops during the night, be sure to let your Observing
Specialist know. Any other unusual artifacts in the images or 
otherwise odd behavior should also be noted immediately.

John Barentine/Mark Klaene/Ed Turner
Mark Klaene
Deputy Site Manager					voice: 505 437 6822
PO Box 59 						fax:   505 434 5555
Sunspot NM 88349-0059					web:

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 733 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to